OpenStudy (anonymous):

Please help? Thanks and show how to so I know for later on. Thank you! :D Keisha's parents want to save $20,000 in her college savings account over the next 15 years. They have $8,000 to use as an initial deposit. What simple annual interest rate do they need to meet their goal? Round your answer to the nearest tenth. A.) 12% B.) 10% C.) 2.7% D.) 2.5%

OpenStudy (anonymous):

"Simple interest is calculated on the original principal only" interest rate = x , after 1 year they gain 8000x , and they do this for 15 years so: 15(8000x) = 20000 120000x = 20000 x = 20000/120000 = 2/12 =1/6 which is 16.7% .. unless i have made an error somewhere or misinterpreted your question..

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

Do I still need to round? 16.7 isn't a choice?

OpenStudy (anonymous):

that is why im confused, ill ask on chat for someone to check my working

OpenStudy (anonymous):

Okay thank you for all your help :D

OpenStudy (cwrw238):

i've checked out eigens working - looks fine to me